OpenStudy (anonymous):

The period of a physical pendulum is measured first at sea level, where the pendulum completes 100 cycles in 1550 s, and then on a mountaintop, where it comepletes 100 cycles in 1551 s. Assuming that the change in period is caused only by the variation in g with altitude, find the height of the mountain.

OpenStudy (irishboy123):

go find the equation for the period of pendulum and you will see T is proportional to 1/ root (g) or T = A / root (g) where A = const. you can also use Newton's grav equation F = GMm/r^2, meaning that g is inverseley proportional to distance from earth's centre or g = B / r^2 where B = const. combining these two gives you a very simple relationship between T and r. remembering of course that r = distance from earth's centre and not height above sea level.

OpenStudy (anonymous):

this on is hard ye? I will have to find some relation between force of gravity and length of pendulum? @IrishBoy123

OpenStudy (irishboy123):

yes, see above

OpenStudy (anonymous):

the equation for pendulum is\[T=2\pi*\sqrt{l/g}\]? \[g=B/r^2. Solve.for.r=> r=\sqrt{B/g}\] and then B is amplitude?

OpenStudy (anonymous):

or is B the period? 15.5 s?

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (irishboy123):

A and B are just constants that should cancel out when you do the arithmetic.....

OpenStudy (anonymous):

could you write the equations in equations?

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (irishboy123):

\[T = \frac{A}{\sqrt{g}}\] \[g = \frac{B}{r^{2}}\] \[T = \frac{A}{\sqrt{{\frac{B}{r^{2}}}}} = Cr, where C = constant \] Thus, T is directly proportional to r, I think. do you? if so we cany say T1/T2 = r1/r2 and solve that way, assuming you agree....

OpenStudy (anonymous):

I did it like this.. \[T=2\pi \sqrt{l/g}\] solve for l \[l=g*T^2/4\pi^2\] \[l_1=l_2\] so \[g_1*T^2_1/4\pi^2=g_2*T^2_2/4\pi^2\] and then 4pi^2 can be equalised so we have: \[g_1*T^2_1=g_2*T^2_2\] solve for g_2 since g1 is at sea level. \[g_2=9.82m/s^2*(15.5s)^2/(15.51s)^2 = g_2=9.807m/s^2\] from there I used \[F=G*Mm/r^2=m*g_2\] solve for r: \[r=\sqrt{G*M/g_2}\] and the answear from that is minus the earth radius?

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (anonymous):

Or can I put in g2 in the l equation? in that case I get that the length is 3.85 meters?

OpenStudy (irishboy123):

your methodology looks good to me. i just cut out some more of the variables and thus the middle step. i get your answer to be 6374km. and yes, you need to subtract 6371 from that, as that is the ave radius from Wiki (reliable?!?!). meaning the mountain is 3 km high. my method says 15.51/15.50 = r/6371 r = 6375.11 mountain = c 4 km. both methods get you in the ballpark and demonstrate an understanding of the physics. this is not an inclined plane problem where the answer is exactly 2.5 or π!!

OpenStudy (anonymous):

I get my answear to 6372 km :/ maybe its calculater wrong...

OpenStudy (anonymous):

so that the mountain is 1km high :S

OpenStudy (irishboy123):

yeah, my method doen't need G or M but it is a bit at the mercy of the radius figure you use. i get 6379 for you method when i use the constants in my physics book. that is probably the best place to look for them. wiki can be all over the place.

OpenStudy (anonymous):

ye true! But what constants for G and M do you have? G=6.67*10^-11? M=5.97*10^24? @IrishBoy123